Maximise 1170x1 + 1110x2

Subject to: 9x1 + 5x2 ≥ 500 7x1 + 9x2 ≥ 300 5x1 + 3x2 ≤ 1500 7x1 + 9x2 ≤ 1900 2x1 + 4x2 ≤ 1000 x1, x2 ≥ 0 -Find graphically the feasible region and the optimal solution


1
Expert's answer
2022-03-07T01:31:09-0500

1170x1 + 1110x2 is maximized as follows = (1170*1)+(1110*2) =3390

The image below shows the relevant graph.